# EgoHand: Ego-centric Hand Pose Estimation and Gesture Recognition with Head-mounted Millimeter-wave Radar and IMUs

### v1

@@ 4 Experiments
4.1 Dataset Acquisition
Dataset was collected under the approval of IRB.
(1) Hardware Configurations. The data collection hardware consists of a millimeter-wave radar, an IMU, and a camera, shown in Fig. 5. These devices were manually synchronized before recording started.
$\bullet$ Millimeter-wave radar. One TI IWR6843ISK radar generates frequency modulated continuous wave (FMCW) signals at frequencies of 60GHz-64GHz. Its transmit power is only 15.85 mW, slightly lower than typical Wi-Fi devices, making it safe for human exposure. The radar has three transmitting antennas and four receiving antennas. The transmit parameters are set to 20 frames per second with 64 chirps per frame. The analog-to-digital converter (ADC) sampling rate was 256, and the raw radar data was obtained by mounting the radar on a TI DCA1000EVM board. The raw radar data is $\in 20t\times 3\times 4\times 256\times 64$, where $t$ is the time in seconds. The range-Doppler and range-angle heatmaps are obtained after the preprocessing process, both of which are $\in 20t\times 256\times 64$.
@@ 4.1 Dataset Acquisition
(2) Dataset diversity. We recruited 10 volunteers with heights ranging from 161 to 182 centimeters and weights between 44 to 73 kilograms. We conducted our experiment across three different scenes and three distinct postures, i.e., sitting, lying, and standing, wherein each volunteer engaged in 8 types of human-computer interaction gestures. These gestures include click, swipe leftward, swipe rightward, swipe upward, swipe downward, swipe to bottom right, zoom in, and zoom out. The latter two are two-handed gestures and the rest are single-handed gestures. The selection of these gestures and postures is inspired by the Apple Vision Pro, such as confirming selection and scrolling up the page, which has demonstrated that this set of gestures can facilitate a wide range of VR interactive functions, as shown in the promotional videos released by Apple [^1] Youtube link and Bilibili link. In each scene and posture, volunteers were instructed to perform each gesture 20 times, with each gesture completed within 2 seconds.
Of the 10 volunteers, two are left-handed, and in performing single-handed gestures, they have been assigned tasks to execute using their left and right hands, respectively. The remaining eight volunteers are right-handed, among whom six have been arranged to carry out interactive gestures using their dominant right hand, while two have been assigned to perform gestures using their non-dominant left hand. Thus, there are $2\times 2+6+2=12$ different settings for the left and right hand executing gestures in total. Table. 1 shows this experimental setting. This setup has been designed for scenarios where a user’s preferred hand is occupied, ensuring that the system can still recognize and respond to gestures executed by the non-preferred hand.
@@
Table 1: We recruited 10 volunteers to perform VR interaction gestures. They were asked to execute gestures by their preferred hand or non-preferred hand. Pre.Hand for Preferred hand; Exe.Hand for Executing Hand; H. for Height; W. for Weight.
(3) Dataset statistics. The dataset contains 5760 (3 scenes $\times$ 12 hand execuring settings $\times$ 8 gestures $\times$ 20 repetitions) gesture samples, including synchronized data of millimeter-wave radar, IMUs, and videos. We uniformly select the 5th, 10th, 15th, and 20th repetitions for test, and the remaining 16 repetitions for training. We utilize Google MediaPipe Hand Landmark SDK to extract 3D hand keypoints from videos as the ground truth for deep network training and evaluation. We observe that the SDK sometimes misses to extract hand keypoints, especially in situations when hands move very fast. If the SDK misses 40% or more frames in a particular video, we discard it. Otherwise, we downsample or upsample the hand keypoints to 30 frames per gesture sample. Although MediaPipe’s 3D positions may lack high precision, this approach significantly simplifies the annotation process. Consequently, we have a dataset that consists of 5206 gesture samples, detailed statistics shown in Table. 2.
@@
Table 2: Dataset samples. Volunteers executed the gestures with postures of sitting in scene #1, lying in in scene #2, and standing at scene #3, respectively.
@@ 4.2 Evaluation Metrics
4.3 In-domain Results
Overall performance. Table. 3 presents the main results of EgoHand  under the data splitting strategy shown in Table. 2. The one-stage approaches directly output the gesture classification results without regressing the hand keypoints. In contrast, EgoHand  outputs hand keypoints as the intermediate features and conduct gesture recognition over the keypoints. We evaluate current mainstream and basic classification networks. Table. 3 demonstrates that EgoHand  outperforms one-stage approaches by a large margin (except EgoHand  (ViT [10])). This result demonstrates that the novel end-to-end Transformer architecture proposed in our first stage is highly effective. A more intriguing finding is that, despite not achieving particularly good MPJPE in the first stage, intermediate hand keypoint positions significantly boost the accuracy of gesture recognition. This suggests that future work could consider promoting hand keypoint estimation to enhance gesture recognition. In light of the good performance of ResNet50 in classification tasks, this network has been selected as the gesture recognition backbone.
Metrics MPJPE Acc(%) Precision Recall F1
@@ 4.3 In-domain Results
one-stage (ViT) / 59.66 67.12 59.77 60.75
one-stage (ResNet50) / 60.09 61.79 61.19 60.64
one-stage (mGesNet) / 81.34 81.98 81.56 81.64
one-stage (mSeeNet) / 84.60 86.73 86.59 86.60
one-stage (mmGesture) / 68.66 70.01 68.83 68.81
EgoHand (GCN) 81.95 85.36 87.16 87.12 87.09
EgoHand (LSTM) 75.70 89.26 91.56 91.48 91.46
EgoHand (ViT) 91.20 77.95 81.39 80.19 79.28
EgoHand (ResNet50) 72.73 90.80 93.34 93.20 93.19
Table 3: EgoHand outperforms one-stage methods such as mGesNet [24], mSeeNet [26], and mmGesture [40], by a large margin. MPJPE is measured in millimeters(mm).
@@
Table 4: Accuracy, Precision, Recall and F1 for each gesture.
Figure 6: EgoHand  hand pose estimation examples, showing click, swipe downward, zoom in, and swipe leftward, respectively. (Images here are only used for visualization).
@@
PJPEs for fingers. In addition, PJPEs of the thumb finger, index finger, middle finger, ring finger, and little finger are 60.18, 70.10, 76.42, 84.44, and 91.79, respectively. The thumb and index finger exhibit the least prediction errors, whereas the largest errors are found in the ring and little fingers. This is because, from the perspective of a mmWave radar worn on the head, the thumb and index finger are the most discernible, while the other fingers tend to be frequently obscured by the thumb and index finger.
@@
Visualization. Fig. 6 illustrates the visualization of hand keypoint predictions by the EgoHand  for various gestures, showcasing different volunteers performing gestures such as click, swipe downward, zoom in, and swipe leftward. Each set of gestures is depicted in three rows: the first row presents the video frames captured during the gestures, the second row shows the hand keypoints derived from the video frames using the Google SDK, and the third row displays the predicted hand keypoints output by EgoHand. It can be seen that EgoHand  can automatically detect the number of hands, whether they are left or right hands, as well as the movement of hand joints throughout the gesture executing process, which effectively alleviates the difficulty of subsequent gesture recognition.
Gesture recognition. Table. 4 shows the Accuracy, Precision, Recall and F1 Score for each gesture. We can see that the accuracy of single-hand gestures is generally above 90%, whereas the accuracy of two-hand gestures is slightly lower but still remains above 80%. The most dominant classification errors are the mutual misclassifications between the two-hand gestures of zoom in and zoom out. This is because, during the data collection, volunteers have no strict guidance on which diagonal direction should be used for zoom in and zoom out, introducing more variability compared to single-hand gestures. Besides, The MPJPEs of zoom in and zoom out in the first stage are generally higher than the single-hand gestures, leading to suboptimal performance in the subsequent gesture classification.
Modality ablation study. Table. 5 shows the experiments relating to the impact of IMU data in EgoHand. The experiments employed strategies using only millimeter-wave radar data, only IMU data, and both data streams. Results indicate that using only IMU data to estimate gestures results in weakest performance in terms of both mean MPJPE and gesture recognition accuracy. Relying only on mmWave radar data also leads to a decrease in metrics compared to the both using millimeter-wave radar and IMU data. This experiment demonstrates that data captured by IMUs can effectively compensate for human body or head movements, thereby enhancing the accuracy of hand keypoint estimation as well as gesture recognition.
4.4 Cross-domain Evaluation
Cross-person. We utilize two cross-person validation approaches to evaluate the performance of EgoHand  in gesture recognition on unseen users. (1) We use dataset from No.1 to No.8 listed in Table. 1 as the training set to train EgoHand, and test the trained EgoHand  with dataset from No.9 to No.12. (2) We select one of 12 rows from Table. 1 for testing and use the remaining rows for training; and we iteratively repeat this process 12 times and ultimately compute the average value. Table. 6 shows that EgoHand  maintains a certain level of recognition capability for unseen persons, achieving a gesture recognition accuracy of 63.11 and 61.96% in the above two training/test splitting strategy, respectively. Moreover, the gesture recognition accuracy significantly improves by 10%-20% with one-shot or two-shot fine-tuning, using only one or two samples from the target person/s to fine-tune EgoHand.
Input MPJPE Acc(%) Precision Recall F1
IMUs 120.18 68.80 70.13 69.80 69.36
mmWave 96.42 86.04 88.70 88.55 88.54
Fusion 72.73 90.80 93.34 93.20 93.19
Table 5: With mmWave radar and IMUs, EgoHand  performs the best. MPJPE is measured in millimeters(mm).
@@ 4.4 Cross-domain Evaluation
Table 6: Gesture recognition tests on unseen volunteers. EgoHand  maintains a certain level of recognition capability for unseern persons, and can be improved by 10%-20% with one-shot or two-shot fine-tuning.
Figure 7: Gesture recognition accuracy under leave-one-person-out setting.
@@
sitting in scene #1 (in-domain result dataset split) / 87.60
lying in scene #2 (in-domain result dataset split) / 87.96
standing in scene #3 (in-domain result dataset split) / 95.92
@@
Table 7: Gesture recognition tests on unseen posture/scene. EgoHand  maintains a certain level of recognition capability for unseen posture/scene, and can be improved by 20% with one-shot or two-shot fine-tuning.
Fig. 7 demonstrates the leave-one-person-out results. It also shows that EgoHand  maintains a certain level of recognition capability and can be improved further by few-shot fine-tuning. There is one exception observed with No. 7, which exhibits a significantly lower accuracy rate. Upon reviewing the video footage, we find that this particular volunteer performs the gestures with smaller and more casual movements, deviating considerably from the patterns exhibited by others. Consequently, during actual system deployment, what we need is to implement a pre-use calibration protocol, requiring users to calibrate each gesture twice or more before commencing usage.
Cross-posture and Cross-scene. We use data from two postures (or scene) in Table. 2 to train EgoHand and tests the trained EgoHand on data of the last postures (or scene). Similar to the cross-person performance, EgoHand  maintains a certain level of recognition capability for unseen postures/scenes, achieving a gesture recognition accuracy of 54.35%, 51.42%, and 51.41% in three cross-posture evaluation, respectively. Moreover, the gesture recognition accuracy significantly improves by 20% with one-shot or two-shot fine-tuning. This experiment further substantiates the efficiency of pre-use calibration strategy discussed in the above cross-person evaluation.
@@
In-domain results accuracy of left hand / 90.96
@@
Table 8: Gesture recognition tests on unseen hands. EgoHand  maintains a weak level of recognition capability for unseern hands, and can be largely improved by 30%-50% with one-shot or two-shot fine-tuning.
Cross-hand. Table. 8 shows that EgoHand  achieves that the accuracy for each hand exceeds 90% if following the data splitting strategy in in-domain results section. We also conduct cross-hand evaluation, training EgoHand  with data of right/left executing hand in Table. 1, and test the trained EgoHand  with data of left/right executing hand, respectively. It is noteworthy that this training strategy also encompasses cross-person challenges. For example, as shown in Table. 1, instances from No.6 to No.11 introduce cross-person challenges when the left-hand execution data is used for training and the right-hand execution data is employed for testing. Thus, this is a highly challenging data splitting scheme, with Table. 8 indicating that EgoHand  struggles greatly under such split. However, upon fine-tuning with just one to two samples, gesture recognition accuracy sees a remarkable improvement of 30-50%, up to 73.53% and 77.18%. This indicates that during the training of EgoHand, we should simultaneously provide samples from both the left and right hands.

### v2

@@ 4 Experiments
4.1 Data Acquisition
Data was collected under the approval of the IRB.
(1) Hardware Configurations. The data collection hardware consists of a millimeter-wave radar, an IMU, and a camera, shown in Fig. 5. The millimeter-wave radar and IMU device are mounted on a plastic bracket and secured to the head using an adjustable strap. A tripod is used to position the camera for hand motion capture and hand pose labeling. All three devices are controlled and synchronized via a desktop computer, ensuring precise temporal synchronization during data recording.
$\bullet$ Millimeter-wave radar. One TI IWR6843ISK radar generates frequency modulated continuous wave (FMCW) signals at frequencies of 60GHz-64GHz. Its power consumption is only 1.75W, slightly lower than typical Wi-Fi devices, making it safe for human exposure. The radar has three transmitting antennas and four receiving antennas. The transmit parameters are set to 20 frames per second with 64 chirps per frame. The analog-to-digital converter (ADC) sampling rate is 256, and the raw radar data is obtained by mounting the radar on a TI DCA1000EVM board. The raw radar data is $\in 20t\times 3\times 4\times 256\times 64$, where $t$ is the time in seconds. The range-Doppler and range-angle heatmaps are obtained after the preprocessing process described in Sec. 3.1, both of which are $\in 20t\times 256\times 64$. These heatmaps are then combined along the distance dimension to obtain model inputs $\in 20t\times 256\times 128$.
@@ 4.1 Data Acquisition
(2) Dataset diversity. We recruited 10 volunteers, whose heights ranged from 161 to 182 centimeters and weights from 44 to 73 kilograms. The experiment was conducted across three different environments and under three distinct postures: sitting, lying, and standing. In each condition, the participants performed eight types of human-computer interaction gestures: click, swipe left, swipe right, swipe up, swipe down, swipe to the bottom-right, zoom in, and zoom out. Among these, the last two are two-handed gestures, while the rest are performed with a single hand. The selection of gestures and postures was inspired by the Apple Vision Pro, which employs similar interactions—such as confirming selections or scrolling through content—for a broad range of VR functionalities. In every scene and posture, each volunteer was asked to perform each gesture 20 times, with each gesture completed within a 2-second duration.
Among the 10 volunteers, two are left-handed, and in performing single-handed gestures, they were instructed to perform gestures using their left and right hands, respectively. The remaining eight volunteers are right-handed; six of them performed gestures with their dominant right hand, while the other two were assigned to use their non-dominant left hand. Consequently, there are $2\times 2+6+2=12$ different settings for the left and right hand executing gestures in total. Table. II summarizes this experimental configuration. This setup has been designed for scenarios where a user’s preferred hand is occupied, ensuring that the system can still recognize and respond to gestures executed by the non-preferred hand.
TABLE II: We recruited 10 volunteers to perform VR interaction gestures. They were asked to execute gestures by their preferred hand or non-preferred hand. Pre.Hand for Preferred hand; Exe.Hand for Executing Hand; H. for Height; W. for Weight.
@@
(3) Dataset statistics. The dataset comprises a total of 5,760 gesture samples, collected across 3 scenes, 12 hand-executing settings, 8 gesture types, and 20 repetitions per gesture (3×12×8×20). Each sample includes synchronized data from millimeter-wave radar, IMUs, and video recordings. For data partitioning, we uniformly select the 5th, 10th, 15th, and 20th repetitions for testing, and use the remaining 16 repetitions for training. To provide supervision for deep network training and evaluation, we employ the Google MediaPipe Hand Landmark SDK to extract 3D hand keypoints from the video data. However, we observe that the SDK occasionally fails to detect keypoints, particularly when the hand moves rapidly. If more than 40% of frames in a video are missing keypoints, the corresponding sample is discarded. Otherwise, the extracted keypoints are temporally downsampled or upsampled to 30 frames per gesture sample to maintain consistency. Although MediaPipe’s 3D keypoints may not offer high-precision tracking, this automated approach significantly reduces the annotation burden. After filtering out invalid samples, the final dataset contains 5,206 gesture instances. Detailed statistics are presented in Table III.
TABLE III: Dataset samples. Volunteers executed the gestures with postures of sitting in scene #1, lying in in scene #2, and standing at scene #3, respectively.
@@ 4.2 Evaluation Metrics
TABLE IV: mmEgoHand outperforms one-stage methods such as mGesNet [7], mSeeNet [9], and mmGesture [31], by a large margin. MPJPE is measured in millimeters(mm). mm4Arm [27] is a single-hand pose estimation method, trained and evaluated solely on single-hand data. Compared to simultaneous detection of both single-hand and two-hand poses, this is a considerably simpler task. Nevertheless, our method significantly outperforms mm4Arm despite addressing the more challenging setting.
[…]
mmEgoHand (mmWave+IMU) 72.73mm / / / /
mmEgoHand (mmWave) 96.42mm / / / /
mmEgoHand (mmWave+IMU, w/o Context Decoder) 109.02mm / / / /
mm4Arm [27] (mmWave+IMU) 165.19mm / / / /
two-stage (mmEgoHand+ResNet50 [48]) / 90.80 93.34 93.20 93.19
two-stage (mmEgoHand+LSTM) / 89.26 91.56 91.48 91.46
two-stage (mmEgoHand+GCN) / 85.36 87.16 87.12 87.09
two-stage (mmEgoHand+ViT [49]) / 77.95 81.39 80.19 79.28
one-stage (mSeeNet [9]) / 84.60 86.73 86.59 86.60
one-stage (mGesNet [7]) / 81.34 81.98 81.56 81.64
one-stage (mmGesture [31]) / 68.66 70.01 68.83 68.81
@@
one-stage (ResNet50 [48]) / 60.09 61.79 61.19 60.64
one-stage (ViT [49]) / 59.66 67.12 59.77 60.75
4.3 Results
(1) Overall performance.
$\bullet$ IMU works. Table IV presents a comprehensive evaluation of mmEgoHand’s performance in both hand pose estimation and downstream gesture recognition, following the data partitioning strategy detailed in Table III. Our system achieves a Mean Per Joint Position Error (MPJPE) of 72.73mm, demonstrating a 24.6% improvement over mmWave-only approaches (96.42mm MPJPE). This significant enhancement not only confirms the presence of head motion in our data collection process but also validates the effectiveness of our IMU signal fusion strategy for egocentric hand pose estimation.
$\bullet$ Context Decoder works. To further assess the contribution of the Context Decoder, we conducted a controlled experiment by training the model with only the Pose Decoder. In this configuration, the system achieved an MPJPE of 109.92mm. These results confirm that the Context Decoder significantly enhances performance by modeling sequential context across frames.
$\bullet$ SOTA Pose Estimation. Besides, due to the current limitations of mmHand and mm4Arm—namely, their support for single-hand estimation only and their closed-source implementations [25, 26, 27]—we independently reimplemented mm4Arm [27] and trained it exclusively on single-hand data for hand pose estimation, achieving an MPJPE of 165.19mm. This single-hand setting is inherently simpler than our scenario, which simultaneously estimates both single- and two-hand poses. Nevertheless, our method significantly outperforms mm4Arm despite addressing this more challenging task.
$\bullet$ SOTA Gesture Recognition. Furthermore, we use the hand poses estimated by mmEgoHand’s intermediate features and feed them into simple classification models such as ResNet [48], LSTM, GCN, and ViT [49] for gesture recognition, as highlighted in blue in TableIV. All models achieve strong performance under this two-stage pipeline. For instance, using ResNet-50 yields a gesture recognition accuracy of 90.80% and an F1-score of 93.19. These results are 20–30 percentage points higher than directly applying LSTM, ResNet-50, or ViT to raw mmWave and IMU data. Moreover, mmEgoHand  also outperforms carefully designed state-of-the-art methods for mmWave gesture recognition, such as mGesNet [7], mSeeNet [9], and mmGesture [31]. This finding suggests that future research could benefit from incorporating hand pose estimation as an intermediate representation to enhance gesture recognition performance.
TABLE V: Accuracy, Precision, Recall and F1 for each gesture.
@@ 4.3 Results
(2) PJPEs for fingers. In addition, the PJPEs of the thumb, index, middle, ring, and little fingers are 60.18, 70.10, 76.42, 84.44, and 91.79, respectively. The thumb and index fingers exhibit the smallest prediction errors, while the ring and little fingers show the largest. This is because, from the perspective of a head-mounted mmWave radar, the thumb and index fingers are more prominently exposed, whereas the other fingers are often occluded by them, making accurate estimation more challenging. Nevertheless, thanks to the penetration capability of mmWave signals [23], our model is still able to achieve reasonable estimation performance for these occluded fingers.
Figure 6: mmEgoHand  hand pose estimation examples, showing click, swipe downward, zoom in, and swipe leftward, respectively. (Images here are only used for visualization).
@@
(3) Visualization. Fig. 6 illustrates the visualization of hand keypoint predictions by the mmEgoHand  for various gestures, showcasing different volunteers performing gestures such as click, swipe downward, zoom in, and swipe leftward. Each set of gestures is depicted in three rows: the first row presents the video frames captured during the gestures, the second row shows the hand keypoints derived from the video frames using the Google SDK, and the third row displays the predicted hand keypoints output by mmEgoHand. It can be seen that mmEgoHand  can automatically detect the number of hands, whether they are left or right hands, as well as the movement of hand joints throughout the gesture executing process, which effectively alleviates the difficulty of subsequent gesture recognition.
Figure 7: The confusion matrix of gesture recognition.
(4) Downstream gesture recognition. Fig. 7 presents the confusion matrix for recognizing eight human-computer interaction gestures. Each cell in the matrix indicates both the percentage and the absolute number of samples classified into each category. Table V further reports the Accuracy, Precision, Recall, and F1 Score for each gesture. We observe that single-hand gestures generally achieve high accuracy, often exceeding 90%, while two-hand gestures exhibit slightly lower performance, though still above 80%. The most prominent recognition errors occur between the two-hand gestures zoom in and zoom out, which are frequently misclassified with each other. This confusion is largely attributed to the lack of strict instructions given to participants during data collection regarding the specific diagonal direction of movement for these gestures, resulting in greater intra-class variability. Additionally, we observe that the MPJPE for zoom in and zoom out is notably higher compared to single-hand gestures, contributing to reduced accuracy in the downstream gesture classification task.
4.4 Few-shot Gesture Recognition
We further evaluate the generalization ability of mmEgoHand for gesture recognition under few-shot fine-tuning in challenging real-world scenarios, including cross-person, cross-posture, and cross-hand settings. Please note that the primary goal of this evaluation is to demonstrate the baseline generalization capability of our framework. Achieving stronger generalization performance would require incorporating additional techniques such as domain adaptation and domain generalization. To facilitate future research in this direction, we release our dataset publicly to support further exploration of these advanced generalization methods.
(1) Cross-person. We adopt two cross-person validation strategies to evaluate the gesture recognition performance of mmEgoHand  on unseen users. (1) In the first strategy, we use data from No.1 to No.8 (as listed in Table. II) for training, and test the model on data from No.9 to No.12. (2) In the second strategy, we perform leave-one-subject-out evaluation: in each iteration, data from one row in Table. II is used for testing, while data from the remaining rows is used for training. This process is repeated for all 12 rows, and the average performance is reported. As shown in Table VI, mmEgoHand demonstrates a certain degree of generalization ability to unseen users, achieving gesture recognition accuracies of 63.11% and 61.96% under the two strategies, respectively. Furthermore, the accuracy improves significantly—by 10% to 20%—when one-shot or two-shot fine-tuning is applied, using only one or two labeled samples from the target user to adapt the mmEgoHand.
TABLE VI: Gesture recognition tests on unseen volunteers. mmEgoHand  maintains a certain level of recognition capability for unseern persons, and can be improved by 10%-20% with one-shot or two-shot fine-tuning.
@@ 4.4 Few-shot Gesture Recognition
Figure 8: Gesture recognition accuracy in leave-one-person-out setting.
Fig. 8 demonstrates the leave-one-person-out results. It also shows that mmEgoHand  maintains a certain level of recognition capability and can be improved further by few-shot fine-tuning. There is one exception observed with No. 7, which exhibits a significantly lower accuracy rate. Upon reviewing the video footage, we find that this particular volunteer performs the gestures with smaller and more casual movements, deviating considerably from the patterns exhibited by others. Consequently, during actual system deployment, what we need is to implement a pre-use calibration protocol, requiring users to calibrate each gesture twice or more before commencing usage.
(2) Cross-posture and Cross-scene. We use data from two postures (or scene) in Table. III to train mmEgoHand and tests the trained mmEgoHand on data of the last postures (or scene). Similar to the cross-person performance, mmEgoHand  maintains a certain level of recognition capability for unseen postures/scenes, achieving a gesture recognition accuracy of 54.35%, 51.42%, and 51.41% in three cross-posture evaluation, respectively. Moreover, the gesture recognition accuracy significantly improves by 20% with one-shot or two-shot fine-tuning. This experiment further substantiates the efficiency of pre-use calibration strategy discussed in the above cross-person evaluation.
TABLE VII: Gesture recognition tests on unseen posture/scene. mmEgoHand  maintains a certain level of recognition capability for unseen posture/scene, and can be improved by 20% with one-shot or two-shot fine-tuning.
@@
sitting in scene #1 (dataset split in Table III) / 87.60
lying in scene #2 (dataset split in Table III) / 87.96
standing in scene #3 (dataset split in Table III) / 95.92
@@
TABLE VIII: Gesture recognition tests on unseen hands. mmEgoHand  maintains a weak level of recognition capability for unseern hands, and can be largely improved by 30%-50% with one-shot or two-shot fine-tuning.
@@
Table III accuracy of left hand / 90.96
@@
(3) Cross-hand. Table. VIII shows that mmEgoHand  achieves that the accuracy for each hand exceeds 90% if following the data splitting strategy (Table III). We also conduct cross-hand evaluation, training mmEgoHand  with data of right/left executing hand in Table. II, and test the trained mmEgoHand  with data of left/right executing hand, respectively. Note that this training strategy also encompasses cross-person challenges. For example, as shown in Table. II, instances from No.6 to No.11 introduce cross-person challenges when the left-hand execution data is used for training and the right-hand execution data is employed for testing. Thus, this is a highly challenging data splitting scheme. However, upon fine-tuning with just one to two samples, gesture recognition accuracy sees a remarkable improvement of 30-50%, up to 73.53% and 77.18%. This indicates that during the training of mmEgoHand, we should simultaneously provide samples from both the left and right hands.